OpenStudy (anonymous):

(2a+5)sqare of trinomial I want real answer

OpenStudy (anonymous):

\[\huge (2a+5)^2 = (2a)^2 + 2 \times 2a \times 5 + 5^2\] \[\huge = 4a^2 + 20a + 25\] Here we can see that the expansion of (2a+5)^2 gives 3 terms, hence (2a+5)^2 is a trinomial.

OpenStudy (anonymous):

\(\Large \color{red}{use\ Binomial\ Theorem}\ \color{blue}{(a+b)^2=a^2+2ab+b^2}\)